OpenStudy (anonymous):

Please show that (I - 2A) equals its own inverse if A is indempotent--matrices

OpenStudy (anonymous):

\[ MM^{-1}=I \]So you need to show \[ (I-2A)(I-2A)=I \]

OpenStudy (anonymous):

\[ \begin{split} (I−2A)(I−2A) &=I(I−2A)−2A(I−2A)\\ &=II-2A-2A+(-2A)(-2A)\\ &=I-4A+4(AA) \\ &= I-4A+4A\\ &=I \end{split} \]

OpenStudy (anonymous):

Medals. All of them.

OpenStudy (anonymous):

thankx